. A block slides down a ramp that has a length of 11.4 m. The block travels at a speed of 5.43 m/s at the bottom of the ramp and increased in temperature from 22.1°C to 22.3°C. The block is made of copper and has a specific heat capacity of 385 J/kg°C. What is the angle of the ramp?


1
Expert's answer
2021-04-19T17:10:51-0400

"m\vec{g}+\vec{N}+\vec {F_{fr}}=m\vec a,"

"ma=mgsin\alpha-\mu mg cos\alpha,"

"\mu=\frac{gsin\alpha-a}{gcos\alpha},"

"v=at,~l=\frac{at^2}{2}\implies a=\frac{v^2}{2l},"

"A_{fr}=Q,"

"\mu mg\cos \alpha l=cm\Delta t,"

"(gsin\alpha-\frac{v^2}{2l})l=c\Delta t,"

"sin\alpha=\frac{v^2+2c\Delta t}{2gl}=0.82,"

"\alpha=55° ."